OpenStudy (anonymous):

Is f(x)=1/x uniformly continuous then x is in (0,1)

OpenStudy (anonymous):

are you asking if this statement is true?

OpenStudy (anonymous):

Yes, but I need a proof accompanying the statement, of course.

OpenStudy (anonymous):

(0,1) is an open interval. \[x \neq 0\]

OpenStudy (anonymous):

well then its true =)

OpenStudy (anonymous):

And why is that?

OpenStudy (anonymous):

cuz f(x) is defined on the open interval (0,1)

OpenStudy (anonymous):

Well, I know that if a function is continious then it will be uniformly continious in any CLOSED interval. I haven't heard anything about such thing being true for open intervals

OpenStudy (anonymous):

The function is not uniformly continuous for any closed interval. If the interval included 0, it would be undefined.

OpenStudy (anonymous):

I was talking about functions in general. Of course I meant the intervals in which the function is defined

OpenStudy (anonymous):

I think that it's not uniformly continuous so I am trying to prove that there exists an epsilon>0 for all deltas>0 such that there exist x,y such that \[|f(x)-f(y)|\geq \epsilon \] if \[|x-y|<\delta\] I started by choosin epsilon=1. Then: \[|\frac{ 1 }{ x } - \frac{ 1 }{ y }| = \frac{ |x-y| }{ xy }\] Now, I think, I need to choose such values of x and y expressed through delta such that the equality above would be greater or equal than one, but I am having trouble thinking of such values. Is my approach any good?